1. 1) luciferin+O2 ⇌ oxyluciferin+light

2) ATP⇌AMP+PPi ΔG=−31.6 kJ/mol

If the overall ΔG of the coupled reaction is -1.21 kJ/mol , what is the equilibrium constant, K, of the first reaction at 11C? Round your answer to 3 significant figures.

2. When methanol (CH3OH) is combusted, such as when in a gasoline blend, the following reaction occurs:

2CH3OH(l)+3O2(g)→2CO2(g)+4H2O(g)

Based on the standard free energies of formation given, what is the standard free energy change for this reaction? Report the answer in kJ/mol rxn rounded to the ones place.

Substance ΔHf° (kJ/mol) ΔGf° (kJ/mol) S° (J/mol K)

CH3OH(l)

−239.2 −166.6 126.8

O2(g)

0.0 0.0 205.2

CO2(g)

−393.5 −394.4 213.8

H2O(l)

−285.8 −237.1 70.0
Answer #1

(1) шаРЕПАN то — ATP = AMP + PP; wLIFцо , чҷи)ься, °C = 284, 15). D = - 31.667 more. allen Du in posinue the REACTION IS NON-(2) 2 CHOH + 30, (g) 2 co₂ + 4H2O 23 Du f H Chon= -166.665/mor (for aromole). 72. Con= 24-1866Timo =-333.2 kg/mol b) 02 (6)